# Smart Data Portfolios:

### v1

@@ 4.1 Informational Return
Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ denote the available data categories, and let $w=(w_{1},\ldots,w_{n})$ be a nonnegative weight vector with $\sum_{i}w_{i}=1$. Training a model from class $\mathcal{M}$ on the mixture induced by $w$ yields a task-specific performance metric $M(w;\mathcal{M})$, evaluated on an institutionally approved validation procedure. Depending on the application, this may correspond to AUC, cross-entropy loss, RMSE, calibration error, or other sector-appropriate performance indicators.
@@
In practice, institutions estimate $\mu_{p}(w)$ using standard validation protocols already embedded in their ML pipelines, such as cross-validation, temporal splits, or regulator-approved holdout procedures. The SDP framework does not prescribe a particular metric or estimation method, requiring only that informational return be computed consistently and documented in a form suitable for regulatory review.
This treatment mirrors modern portfolio theory, where the efficient frontier is defined relative to an institution’s return model and estimation choices rather than as an absolute property of the assets themselves (Markowitz1952). Accordingly, different institutions may trace different governance-efficient frontiers even under identical regulatory constraints, reflecting heterogeneous modeling strategies rather than governance failures.
@@ 4.2 Governance-Adjusted Risk
Governance-adjusted risk is a composite measure that captures expected governance burdens. These burdens typically arise from fairness dispersion, documentation and provenance defects, or operational instability. Each of these dimensions is explicitly recognized in contemporary AI governance regimes, including the EU Artificial Intelligence Act (EUAIAct2024), the OECD AI Principles (OECD2019), and the NIST AI Risk Management Framework (NIST2023). To operationalize these concerns at the input level, we decompose governance-adjusted risk into three components:
@@
    Fairness Dispersion ($F(w;\mathcal{M})$). Variation in model performance across relevant subgroups, measured through gaps in error rates, acceptance rates, or domain specific fairness indicators (Barocas2016; Hardt2016). As with informational return, this is evaluated on the trained model.
@@
    Provenance and Quality Defect Score ($P(w)$). Documentation and lineage risks such as missing consent metadata, uncertain licensing, or high label noise (Gebru2021). This component is model agnostic, reflecting properties of the underlying data lineage rather than model behavior.
@@
    Robustness Volatility ($R(w;\mathcal{M})$). Instability under benign distribution shifts, estimated from out-of-time or out-of-subset tests (Recht2019; Taori2020). One operational definition is
@@
For comparability across institutions and to eliminate scale ambiguity, each component is mapped to a common unit interval using regulator-defined transformations. As a result, $F(w;\mathcal{M}),P(w),R(w;\mathcal{M})\in[0,1]$, and the relative magnitudes of the policy parameters $(\alpha,\beta,\gamma)$ reflect regulatory priorities rather than arbitrary metric scales.
@@
The composite quantity $\sigma_{p}(w)$ treats governance burdens as an average-case measure. In high-stakes settings, however, regulators are primarily concerned with tail exposure, which refers to rare but severe fairness, privacy, or provenance failures. Consequently, a natural extension replaces $\sigma_{p}(w)$ with its Conditional Value-at-Risk (CVaR) analogue (Rockafellar2000).
@@
Following Rockafellar2000, CVaR admits a convex reformulation via auxiliary variables, enabling implementation through linear programming. Regulators may therefore replace the policy risk cap with $\sigma^{\text{CVaR}}_{p}(w)\leq\bar{\sigma}$ when tail control is required. This refinement provides a more advanced risk measure aligned with contemporary supervisory practice in financial regulation.
@@ 4.3 Governance-Efficient Frontier
    For each candidate, train or fine tune the model class $\mathcal{M}$ on the induced mixture and evaluate the validation metric and governance components $F(w^{(k)};\mathcal{M})$, $P(w^{(k)})$, and $R(w^{(k)};\mathcal{M})$.
@@
Because informational return and governance-adjusted risk are not statistically dual quantities, the shape of the frontier in $(\sigma_{p}(w),\mu_{p}(w))$ space is empirical. Existing studies, summarized in Section 2, suggests that in many regulated consumer-facing settings the frontier is approximately concave or inverse-U shaped: marginal informational gains saturate once core categories are included (Recht2019; Taori2020), while fairness, privacy, and provenance risks typically rise with additional behavioral or weakly curated inputs (Mehrabi2021; Sambasivan2021). A detailed empirical analysis of frontier shapes and associated confidence intervals is left for future work.
@@ 4.4 Portfolio optimization and regulation
Once the governance-efficient frontier is established, portfolio optimization proceeds by selecting the data allocation that maximizes validated performance subject to regulator-defined governance constraints. The central regulatory instrument is the policy risk cap, which sets an upper bound on aggregate governance exposure while preserving institutional discretion over model architectures, training procedures, and validation strategies.
@@
In practice, optimization is carried out over a constrained feasible region defined jointly by the policy risk cap and additional category-level governance instruments. Because neighboring portfolios correspond to nearby weight vectors, warm starts and incremental fine-tuning substantially reduce computational cost. For high-dimensional or expensive training regimes, Bayesian optimization with Gaussian Process surrogates provides an efficient exploration strategy (Snoek2012; Frazier2018).
@@
$\sum_{i\in H}w_{i}=0\qquad\text{or}\qquad\sum_{i\in H}w_{i}\leq\beta.$ (7)
@@ 4.5 Attribution and interpretability
Although the frontier uses the aggregate quantities $(\mu_{p}(w),\sigma_{p}(w))$, institutions may require transparency regarding the contribution of individual data categories. Because $M(w)$ is generally nonlinear in $w$, additive decompositions of the form $\mu_{p}\approx\sum_{i}w_{i}\mu_{i}$ lack theoretical justification. As a result, a cooperative game theoretic formulation is required.
Let $\mathcal{N}=\{1,\ldots,n\}$ be the set of categories and let $M(S)$ denote performance when training only on categories $S\subseteq\mathcal{N}$. The Shapley value for category $i$ (Shapley1953; Strumbelj2014; Ghorbani2019) is given as:
@@
Shapley values provide transparent attribution without relying on model specific heuristics. Importantly, they do not influence frontier construction, which depends solely on aggregate portfolio level quantities. They instead serve a justificatory role, supporting explanation and contestation of portfolio choices without altering governance constraints or optimization outcomes.
While exact computation is exponential in $n$, practical estimators for $\phi_{i}$ based on Monte Carlo sampling, influence functions, or gradient based methods are well established (Koh2017). For portfolios with a moderate number of categories, sampled permutations on the order of a few hundred to a few thousand often yield stable estimates in empirical studies (Jia2019).
@@ 5 Policy and Implementation
The Smart Data Portfolio (SDP) framework is designed as an input-level governance architecture rather than a model-level control mechanism. Its central policy contribution is to make data-selection decisions explicit, measurable, and auditable, while preserving institutional discretion over model choice, validation procedures, and optimization routines. By operating at the level of data allocation, SDPs translate high-level governance objectives into enforceable constraints without prescribing technical implementations, consistent with emerging standards such as ISO/IEC 42001 (ISO42001). This practical focus is especially important for agentic AI systems, whose adaptive and multi-step decision processes can render output-level explanations unstable over time.
@@
The framework is most immediately actionable for three groups. First, sector regulators and supervisory agencies face the task of translating broad statutory mandates into auditable, operational rules for high-risk AI systems. Second, specialised AI-governance consultancies increasingly require quantitative, defensible artifacts to support compliance, auditing, and advisory services across jurisdictions. Third, in-house risk, compliance, and data-governance teams in data-intensive firms bear direct accountability for data-selection decisions and must justify those choices to regulators, auditors, and senior management.
These actors share three structural characteristics: direct responsibility for data governance outcomes, authority to impose portfolio constraints, and a practical need for model-agnostic tooling. By framing data allocation as a portfolio-optimization problem, SDPs provide these stakeholders with a common vocabulary and a credible pathway for implementing transparency, fairness, privacy, and robustness requirements across heterogeneous AI systems.
@@
Policy implementation under the SDP framework proceeds iteratively. Regulators specify governance parameters, including admissible data categories, governance weight bands, and an overall policy risk cap. Institutions then construct candidate data portfolios and evaluate them using the informational and governance-risk metrics defined in Section 4. Within these constraints, institutions retain full discretion over model class, training procedures, and validation strategies, but must justify resulting input allocations in a structured and reproducible manner.
Supervisory oversight occurs through periodic review of submitted portfolios. When constraints are violated or governance thresholds exceeded, regulators may require rebalancing or impose tighter bounds. As data environments evolve through concept drift, the introduction of new data sources, or changes in legal requirements, institutions revisit the portfolio-optimization problem and adjust weights accordingly. This cyclical structure mirrors established practices in financial supervision and model-risk governance, adapted to the specific challenges posed by data-centric AI systems.
@@
The Data Portfolio Statement (DPS) provides a public-facing summary of admissible data categories, applicable governance weight bands, and the principles governing input allocation. The Data Portfolio Card (DPC) serves as a regulatory filing that records realised category-level weights together with validated values of informational return $\mu_{p}(w)$ and governance-adjusted risk $\sigma_{p}(w)$, supported by documentation of data sources, lineage, and validation procedures. The Consumer Portfolio Report (CPR) provides consumer-facing explanations of adverse actions or decisions by referencing the approved input portfolio rather than internal model mechanics.
@@
Existing documentation tools, most notably Model Cards and Datasheets (Mitchell2019; Gebru2021), improve transparency by describing model behavior and individual dataset properties. However, they do not govern how multiple data sources are combined, weighted, or constrained within a training mixture. Smart Data Portfolios operate at this missing allocation layer. While Model Cards summarize model performance and Datasheets document dataset lineage and quality, SDPs provide portfolio-level explanations through standardized reporting artifacts. As AI systems increasingly mediate regulated services, this input-governance layer becomes the primary layer of accountability and explainability.
Table 1 summarises the core policy instruments within the SDP framework. Each instrument functions at the portfolio level and is designed to complement, rather than replace, existing documentation practices.
@@ 6 Sectoral Illustration: Telecom as a Multi-Service AI Platform
Telecommunications operators provide an excellent test case for Smart Data Portfolios because they operate multi-service AI platforms built on heterogeneous, high-volume, and governance-sensitive data ecosystems. Privacy obligations arise under general data-protection regimes such as the GDPR, including data-minimization and purpose-limitation principles, as well as under sector-specific telecom regulations, including customer proprietary network information (CPNI) rules in the United States and parallel constraints under European data-protection and ePrivacy frameworks (FCC2023; EDPB2018). These regulatory obligations align closely with governance concerns emphasized in the academic literature (Acquisti2016; Barocas2016; Wachter2019; Recht2019; Mehrabi2021; Jakubik2024). Yet in telecom practice, these requirements remain largely procedural, offering limited guidance on how heterogeneous data sources should be selected, weighted, and constrained within AI training pipelines.
@@ 6.1 Data categories and governance profiles
A portfolio treatment clarifies these distinctions by mapping each category to contributions in informational return and governance-adjusted risk. Because telecom operators support AI across multiple product lines, a single governance envelope is insufficient. Each AI deployment has a distinct fairness, privacy, documentation, and robustness profile. The Smart Data Portfolio framework assigns each use case its own admissible region and governance-efficient portfolio.
@@ 6.2 Representative AI use cases
We illustrate this heterogeneity through three representative AI use cases with distinct governance requirements: device finance scoring, personalization, and network QoS prediction. These stylized examples demonstrate the mechanics of portfolio construction and the governance logic with no empirical claims about real-world performance.
@@
Telecom operators routinely offer handset installment plans and small credit products, placing these models at the intersection of algorithmic fairness requirements and financial regulation (Barocas2016; Hardt2016). The privacy literature also highlights risks associated with linking behavioral data to financial decisions (Acquisti2016; Wachter2019). Regulators therefore impose tight governance constraints. In portfolio form, which can be expressed as:
@@
We consider five data categories ($i=1,\ldots,5$) normalized to unit weights. Table 2 shows these along with their component risk score $\sigma_{i}$.
@@
Many telecom operators support recommendation services through video, music, social-media bundles, and personalized offers. These applications raise significant privacy and provenance concerns. Behavioral telemetry is often sensitive and disproportionately revealing (Acquisti2016).
@@
Here, governance risk is dominated by privacy sensitivity and robustness volatility, given evidence of instability in content recommendation and multimodal systems under distribution shift (Recht2019; Taori2020). The governance-efficient portfolio therefore emphasizes coarse usage aggregates, network KPIs, and minimally sensitive survey or preference data.
@@
Network operators rely on AI to predict congestion, outages, and quality degradation. These models primarily use engineering signals and operational logs, which carry lower governance risks relative to consumer-level data. Because the data are mainly operational and engineering-driven, governance concerns relate mostly to reliability rather than distributional fairness (Recht2019; Jakubik2024).

### v2

@@ 4.1 Informational Return
Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ denote the available data categories, and let $w=(w_{1},\ldots,w_{n})$ be a nonnegative weight vector with $\sum_{i}w_{i}=1$.
Operationally, the portfolio weight $w_{i}$ corresponds to the logged sampling share allocated to data category $i$ within the training pipeline. Depending on the application, this share may be implemented through record counts, token proportions, minibatch sampling frequencies, or allocated training-compute budgets under the institution’s data-loading policy. Recording these shares renders portfolio composition observable, reproducible, and auditable for supervisory review.
Training a model from class $\mathcal{M}$ on the mixture induced by $w$ yields a task-specific performance metric $M(w;\mathcal{M})$, evaluated under a governance-approved validation procedure. Depending on the application, this may correspond to AUC, cross-entropy loss, RMSE, calibration error, or other sector-appropriate performance indicators.
@@
In practice, institutions estimate $\mu_{p}(w)$ using standard validation protocols already embedded in their ML pipelines, such as cross-validation, temporal splits, or other validation procedures accepted within the applicable governance framework. The SDP framework does not prescribe a particular metric or estimation method, requiring only that informational return be computed consistently and documented in a form suitable for regulatory review.
This treatment mirrors modern portfolio theory, where the efficient frontier is defined relative to an institution’s return model and estimation choices rather than as an absolute property of the assets themselves (Markowitz, 1952). Accordingly, different institutions may trace different governance-efficient frontiers even under identical regulatory constraints, reflecting heterogeneous modeling strategies rather than governance failures.
@@ 4.2 Governance-Adjusted Risk
Governance-adjusted risk is a composite measure that captures expected governance burdens. These burdens typically arise from fairness dispersion, documentation and provenance defects, or operational instability. Each of these dimensions is explicitly recognized in contemporary AI governance regimes, including the EU Artificial Intelligence Act (European Parliament and Council of the European Union, 2024), the OECD AI Principles (OECD, 2019), and the NIST AI Risk Management Framework (National Institute of Standards and Technology, 2023). To operationalize these concerns at the input level, we decompose governance-adjusted risk into three components:
@@
    Fairness Dispersion ($F(w;\mathcal{M})$). Variation in model performance across relevant subgroups, measured through gaps in error rates, acceptance rates, or domain specific fairness indicators (Barocas and Selbst, 2016; Hardt et al., 2016). As with informational return, this is evaluated on the trained model.
@@
    Provenance and Quality Defect Score ($P(w)$). Documentation, provenance, and privacy-governance defects such as missing consent metadata, uncertain licensing, high reidentification exposure due to linkability or granularity, or high label noise (Gebru et al., 2021; Narayanan and Shmatikov, 2008; Rocher et al., 2019).
@@
    Robustness Volatility ($R(w;\mathcal{M})$). Instability under benign distribution shifts, estimated from out-of-time or out-of-subset tests (Recht et al., 2019; Taori et al., 2020). One operational definition is
@@
For comparability across institutions and to eliminate scale ambiguity, each component is mapped to a common unit interval using regulator-defined scoring rules and transformations. As a result, $F(w;\mathcal{M}),P(w),R(w;\mathcal{M})\in[0,1]$, and the relative magnitudes of the policy parameters $(\alpha,\beta,\gamma)$ reflect regulatory priorities rather than arbitrary metric scales.
@@
The composite quantity $\sigma_{p}(w)$ is an average-case measure. In high-stakes settings, regulators may instead require tail-sensitive control of rare but severe governance failures, in which case $\sigma_{p}(w)$ can be replaced by a Conditional Value-at-Risk (CVaR) analogue (Rockafellar and Uryasev, 2000).
@@
CVaR admits standard convex reformulations, enabling practical implementation in constrained optimization settings (Rockafellar and Uryasev, 2000).
@@ 4.3 Governance-Efficient Frontier
    For each candidate, estimate the validation metric and governance components by applying the institution’s standard training and evaluation pipeline to the induced mixture.
@@
Because informational return and governance-adjusted risk are not statistically dual quantities, the shape of the frontier in $(\sigma_{p}(w),\mu_{p}(w))$ space is empirical. Evidence summarized in Section 2 suggests that, in many regulated consumer-facing settings, the frontier may be approximately concave or inverse-U shaped: marginal informational gains saturate once core categories are included (Recht et al., 2019; Taori et al., 2020), while fairness, privacy, and provenance risks typically rise with additional behavioral or weakly curated inputs (Mehrabi et al., 2021; Sambasivan et al., 2021). A detailed empirical analysis of frontier shapes and associated confidence intervals is left for future work.
@@ 4.4 Portfolio optimization and regulation
Once the governance-efficient frontier is established, portfolio optimization proceeds by selecting the data allocation that maximizes validated performance subject to regulator-defined governance constraints. The central regulatory instrument is the policy risk cap, which sets an upper bound on aggregate governance exposure while preserving institutional discretion over model architectures, training procedures, and validation strategies. Operationally, the policy risk cap functions as a compliance boundary that determines which data mixtures are admissible for deployment, while leaving model choice and internal training procedures to institutions.
@@
In practice, optimization is carried out over a constrained feasible region defined jointly by the policy risk cap and additional category-level governance instruments. Because neighboring portfolios correspond to nearby weight vectors, warm starts and incremental fine-tuning substantially reduce computational cost. For high-dimensional or expensive training regimes, Bayesian optimization with Gaussian Process surrogates provides an efficient exploration strategy (Snoek et al., 2012; Shahriari et al., 2016).
@@
$\sum_{i\in H}w_{i}=0\qquad\text{or}\qquad\sum_{i\in H}w_{i}\leq\theta_{H}.$ (7)
@@ 4.5 Attribution and interpretability
Although the frontier uses the aggregate quantities $(\mu_{p}(w),\sigma_{p}(w))$, institutions may require transparency regarding the contribution of individual data categories. Because $M(w)$ is generally nonlinear in $w$, additive decompositions of the form $\mu_{p}\approx\sum_{i}w_{i}\mu_{i}$ lack theoretical justification. As a result, a cooperative game-theoretic formulation is required.
Let $\mathcal{N}=\{1,\ldots,n\}$ be the set of categories and let $M(S)$ denote performance when training only on categories $S\subseteq\mathcal{N}$. The Shapley value for category $i$ (Shapley, 1953; Štrumbelj and Kononenko, 2014; Ghorbani and Zou, 2019) is given as:
@@
Shapley values provide transparent attribution without relying on model-specific heuristics. Importantly, they do not influence frontier construction, which depends solely on aggregate portfolio-level quantities. They instead serve a justificatory role, supporting explanation and contestation of portfolio choices without altering governance constraints or optimization outcomes.
While exact computation is exponential in $n$, practical estimators for $\phi_{i}$ based on Monte Carlo sampling, influence functions, or gradient based methods are well established (Koh and Liang, 2017). For portfolios with a moderate number of categories, sampled permutations on the order of a few hundred to a few thousand often yield stable estimates in empirical studies (Jia et al., 2019).
@@ 5 Policy and Implementation
The Smart Data Portfolio (SDP) framework is designed as an input-level governance architecture rather than a model-level control mechanism. Its central governance contribution is to make data-selection decisions explicit, measurable, and auditable, while preserving institutional discretion over model choice, validation procedures, and optimization routines. By operating at the level of data allocation, SDPs deliver enforceable constraints without prescribing technical implementations. This approach aligns with regulatory guidance emphasizing accountability for data-driven harms (Federal Trade Commission, 2021) as well as emerging standards such as ISO/IEC 42001 (ISO/IEC, 2023). This practical focus is especially important for agentic AI systems, whose adaptive and multi-step decision processes can render output-level explanations unstable over time.
@@
The framework is most immediately actionable for three groups. First, sector regulators and supervisory agencies face the task of translating broad statutory mandates into auditable, operational rules for high-risk AI systems. Second, specialized AI-governance consultancies increasingly require quantitative, defensible artifacts to support compliance, auditing, and advisory services across jurisdictions. Third, in-house risk, compliance, and data-governance teams in data-intensive firms bear direct accountability for data-selection decisions and must justify those choices to regulators, auditors, and senior management.
These actors share three structural characteristics: direct responsibility for data governance outcomes, authority to impose portfolio constraints, and a practical need for model-agnostic tooling. By framing data allocation as a governed portfolio-allocation problem, SDPs provide these stakeholders with a common vocabulary and a credible pathway for implementing transparency, fairness, privacy, and robustness requirements across heterogeneous AI systems.
@@
Implementation under the SDP framework proceeds through an iterative supervisory cycle. Regulators specify governance parameters, including admissible data categories, governance weight bands, and an overall policy risk cap. Institutions then construct candidate data portfolios and evaluate them using the informational and governance-risk metrics defined in Section 4. Within these constraints, institutions retain full discretion over model class, training procedures, and validation strategies, but must justify resulting input allocations in a structured and reproducible manner.
Supervisory oversight occurs through periodic review of submitted portfolios. When constraints are violated or governance thresholds exceeded, regulators may require rebalancing or impose tighter bounds. As data environments evolve through concept drift, the introduction of new data sources, or changes in legal requirements, institutions revisit the portfolio-optimization problem and adjust weights accordingly. This cyclical structure mirrors established practices in financial supervision and model-risk governance, adapted to the specific challenges posed by data-centric AI systems. Figure 2 summarizes the supervisory workflow, emphasizing the iterative loop from constraint-setting to portfolio construction, validation, deployment monitoring, and review-driven rebalancing.
Regulator sets governance constraints Admissible categories, weight bands, policy risk cap, reporting rules Institution builds SDP Map sources to categories; choose portfolio weights $w$ under constraints Validation & risk estimation Estimate $\mu_{p}(w)$ and $\sigma_{p}(w)$ under governance-approved procedures Deployment & monitoring Deploy within the risk cap; monitor drift, incidents, and constraint violations Audit / supervisory review Review submissions (DPS/DPC/CPR); assess compliance and governance risk Feedback / rebalancing(drift, incidents, regulatory updates)
Figure 2: Iterative supervisory workflow for Smart Data Portfolios (SDPs), linking regulatory constraint-setting, institutional portfolio construction, deployment monitoring, and review-driven rebalancing.
@@
The Data Portfolio Statement (DPS) provides a public-facing summary of admissible data categories, applicable governance weight bands, and the principles governing input allocation. The Data Portfolio Card (DPC) serves as a regulatory filing that records realized category-level weights together with validated values of informational return $\mu_{p}(w)$ and governance-adjusted risk $\sigma_{p}(w)$, supported by documentation of data sources, lineage, and validation procedures. The Consumer Portfolio Report (CPR) provides consumer-facing explanations of adverse actions or decisions by referencing the approved input portfolio rather than internal model mechanics.
@@
Existing documentation tools, most notably Model Cards and Datasheets (Mitchell et al., 2019; Gebru et al., 2021), improve transparency by describing model behavior and individual dataset properties. However, they do not govern how multiple data sources are combined, weighted, or constrained within a training mixture. Smart Data Portfolios operate at this missing allocation layer. While Model Cards summarize model performance and Datasheets document dataset lineage and quality, SDPs provide portfolio-level explanations through standardized reporting artifacts. As AI systems increasingly mediate regulated services, this input-governance layer operates as a primary layer of accountability and explainability in regulated deployments.
Table 1 summarizes the core policy instruments within the SDP framework. Each instrument functions at the portfolio level and is designed to complement, rather than replace, existing documentation practices.
@@ 6 Sectoral Illustration: Telecom as a Multi-Service AI Platform
Telecommunications operators provide a useful illustration of Smart Data Portfolios because they operate multi-service AI platforms built on heterogeneous, high-volume, and governance-sensitive data ecosystems. These environments combine consumer-facing decision systems, personalized services, and engineering optimization tasks within shared data infrastructures subject to stringent privacy and accountability requirements, including constraints on profiling and automated decision-making under European data-protection governance (European Data Protection Board, 2018). Despite extensive procedural compliance regimes, telecom operators typically receive limited operational guidance on how heterogeneous data sources should be selected, weighted, and constrained within AI training pipelines. The SDP framework clarifies these decisions by expressing input governance as a portfolio-allocation problem.
@@ 6.1 Data categories and governance profiles
A portfolio-based governance perspective clarifies these distinctions by mapping each category to contributions in informational return and governance-adjusted risk. Because telecom operators support AI across multiple product lines, a single governance envelope is insufficient. Each AI deployment has a distinct fairness, privacy, documentation, and robustness profile. The Smart Data Portfolio framework assigns each use case its own admissible region and governance-efficient portfolio.
@@ 6.2 Representative AI use cases
We illustrate this heterogeneity through three representative AI use cases with distinct governance requirements: device finance scoring, personalization, and network QoS prediction. These stylized examples are intended solely to illustrate how governance constraints reshape admissible data allocations under the SDP framework rather than to represent empirical performance claims.
@@
Telecom operators routinely offer handset installment plans and small credit products, placing these models at the intersection of algorithmic fairness requirements and financial regulation (Barocas and Selbst, 2016; Hardt et al., 2016). The privacy literature also highlights risks associated with linking behavioral data to financial decisions (Acquisti et al., 2016; Wachter and Mittelstadt, 2019). Regulators therefore impose tight governance constraints. In portfolio form, which can be expressed as:
@@
We consider five data categories ($i=1,\ldots,5$) with weights summing to one. Table 2 shows these along with their component risk score $\sigma_{i}$.
@@
Many telecom operators support recommendation services through video, music, social-media bundles, and personalized offers. These applications raise significant privacy and provenance concerns. Behavioral telemetry is often sensitive and disproportionately revealing (Acquisti et al., 2016).
@@
Here, governance risk is dominated by privacy sensitivity and robustness volatility, given evidence of instability in content recommendation and multimodal systems under distribution shift (Recht et al., 2019; Taori et al., 2020). The governance-efficient portfolio therefore emphasizes coarse usage aggregates, network KPIs, and minimally sensitive survey or preference data.
@@
Network operators rely on AI to predict congestion, outages, and quality degradation. These models primarily use engineering signals and operational logs, which carry lower governance risks relative to consumer-level data. Because the data are mainly operational and engineering-driven, governance concerns relate mostly to reliability rather than distributional fairness (Recht et al., 2019; Jakubik et al., 2024).